Susan hikes 6 miles east, 15 miles north, 1 mile west and 3 miles south. What is her displacement?


1
Expert's answer
2021-10-08T15:43:53-0400

Let's first find "x"- and "y"-components of the resultant displacement:

"d_x=6\ mi\cdot cos0^{\circ}+15\ mi\cdot cos90^{\circ}+1\ mi\cdot cos180^{\circ}+3\ mi\cdot cos270^{\circ}=5\ mi,"

"d_y=6\ mi\cdot sin0^{\circ}+15\ mi\cdot sin90^{\circ}+1\ mi\cdot sin180^{\circ}+3\ mi\cdot sin270^{\circ}=12\ mi."

We can find the magnitude of the resultant displacement from the Pythagorean theorem:


"d=\sqrt{d_x^2+d_y^2}=\sqrt{(5\ mi)^2+(12\ mi)^2}=13\ mi."

We can find the direction of the resultant displacement from the geometry:


"\theta=tan^{-1}({\dfrac{d_y}{d_x}}),""\theta=tan^{-1}({\dfrac{12\ mi}{5\ mi}})=67.4^{\circ}\ N\ of\ E."